• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:06:28 on December18,2008

    Entry number 253918

    keyword=day shift summary

    8:45 Started beam study for Hall C;
    9:20 BB HV tripped, reset successful;
    9:35 MCC said they are bringing back the beam;
    9:45 No beam so far; Target moved to BeO as requested by MCC;
    9:55 Check beam spot with short BeO run 4579, looks good;
    10:00 resume production on He3;
    11:00 DAQ crash. Some gentlemen from the back room fixed it;

    Run 4578 BCM u3 [ 0]:0.04055
    Run 4579 short BeO run to check spot
    Run 4580 no end of run entry
    Run 4581 BCM u3 [ 0]:0.04073
    Run 4582 BCM u3 [ 0]:0.04055
    Run 4583 BCM u3 [ 0]:0.04060
    Run 4584 BCM u3 [ 0]:0.04057
    Run 4585 continue to swing shift